OpenStudy (anonymous):

Simplify completely quantity x squared minus 100 over quantity x squared plus 2 x minus 120 divided quantity 6 x plus 60 over quantity x minus 2

OpenStudy (tkhunny):

Didn't we already do this? \(\dfrac{x^{2} - 100}{x^{2} + 2x - 120} / \dfrac{6x+60}{x-2}\) Reciprocal and Multiply. \(\dfrac{x^{2} - 100}{x^{2} + 2x - 120}\cdot\dfrac{x-2}{6x+60}\) Factor until everything is factored.